OpenStudy (anonymous):

graph and label the intercepts. y=x^3+2. I know the y intercept is 2

OpenStudy (anonymous):

x intercept solve \[0=x^3+2\] \[x^3=-2\] \[x=\sqrt[3]{-2}\]

OpenStudy (anonymous):

looks just like the graph of \[y=x^3\] only raised up two units

OpenStudy (anonymous):

okay, thats what i had i just wanted to be sure it was right, it just looked weird and i wasn't positive about it having the negative under the radical

OpenStudy (anonymous):

oh but it is a "cubed root" not a square root

OpenStudy (anonymous):

a perfectly good real number. for example \[\sqrt[3]{-8}=-2\]

OpenStudy (anonymous):

only even roots cant have negatives right?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

the answer is= spam!!!

OpenStudy (anonymous):

if you're not going to be helpful, stay off my questions!